A market consists of a single product Apple and two individuals A,B and C. A's demand for apple is given by Q=(16-4P), that of B's is Q =(20-2P) and that of C's is Q=(12-2P). If the market supply equation is Q=4P, then the equilibrium price and equilibrium quantity are?



1
Expert's answer
2021-05-31T15:06:32-0400

Total market demand = demand by A + demand by B + demand by C

Demand A= Q= 16- 4P

Demand B = Q= 20- 2P

Demand C= Q = 12- 2P

Total demand = (16-4P) + (20-2P) +( 12- 2P)

= 48 - 8P = Q

Total market supply Q = 4P

To get the market equilibrium price, we equate the total market demand to the total market supply as follows:

48-8 P= 4P

P = 4

Market equilibrium price= 4

Substituting P by 4 in the market demand equation,

48- 8P = Q

48 - 8(4) =Q

Q= 16

Market equilibrium quantity= 16 units